# Zewei Zhou Tianhui Cai11footnotemark: 1 Seth Z. Zhao  Yun Zhang  Zhiyu Huang Bolei Zhou  Jiaqi Ma

### v1

@@ E.2 Evaluation Metrics
nuPlan (NAVSIM). PDMS is the official benchmark metric in NAVSIM on the nuPlan dataset, with its formulation provided in Eq. S3.
@@
CARLA. We evaluate driving performance in closed-loop testing using four key metrics: Driving Score, Success Rate, Efficiency, and Comfortness following the Bench2Drive benchmark [51]. Driving Score reflects the overall route completion penalized by infraction severity. Success Rate measures the percentage of routes completed without major infractions within the time limit. Efficiency quantifies how well the ego vehicle maintains a reasonable speed relative to surrounding traffic, computed as the average speed ratio over multiple checkpoints. Comfortness assesses the smoothness of the trajectory based on physical signals such as acceleration, jerk, and yaw rate, using expert-derived thresholds.
@@ E.3 Model Training
Our primary base model is trained on the mixed training set of the nuPlan and nuScenes datasets. The ego state contains the vehicle’s current velocity and acceleration. For the Waymo end-to-end driving test, we further fine-tune this base model using the Waymo E2E driving dataset. For the model trained on Waymo, the ego vehicle’s state encompasses current acceleration and a 4-second history of vehicle positions and velocities. For closed-loop simulation testing, we train a separate model using data preprocessed from the CARLA Garage dataset and DriveLM-CARLA annotations. This model is trained with single-view inputs instead of the multi-view setup used in the primary model, and we use a large resolution for the input images (with $28\times 28\times 384$ pixels).
@@ E.4 Inference
We employ a stochastic generation strategy using top-p and top-k sampling to generate reasoning and planning outputs. Specifically, we set the temperature to $0.8$, top-k to $20$, and top-p to $0.2$, promoting diverse yet coherent generations. Higher sampling diversity supports slow thinking modes, enabling the model to produce deeper and more elaborate reasoning chains. In contrast, more deterministic settings align with fast thinking, yielding concise and direct responses. This configuration strikes a balance between exploration and reliability, facilitating the generation of CoT reasoning.
@@
ST-P3 [14] 1.33 2.11 2.90 2.11 0.23 0.62 1.27 0.71 - - - - - - - -
VAD [66] 0.17 0.34 0.60 0.37 0.07 0.10 0.24 0.14 - - - - - - - -
UniAD [65] 0.44 0.67 0.96 0.69 0.04 0.08 0.23 0.12 0.48 0.96 1.65 1.03 0.05 0.17 0.71 0.31
EMMA [37] 0.14 0.29 0.54 0.32 - - - - - - - - - - - -
OpenEMMA [44] 1.45 3.21 3.76 2.81 - - - - - - - - - - - -
OpenDriveVLA-3B [34] 0.14 0.30 0.55 0.33 0.02 0.07 0.22 0.10 0.19 0.58 1.24 0.67 0.02 0.18 0.70 0.30
AutoVLA (action only) 0.27 0.42 0.61 0.43 0.15 0.16 0.25 0.19 0.32 0.67 1.12 0.70 0.14 0.28 0.50 0.31
AutoVLA (w/ CoT) 0.25 0.46 0.73 0.48 0.07 0.07 0.26 0.13 0.33 0.81 1.45 0.86 0.08 0.11 0.85 0.35
@@ Appendix F Additional Results
We evaluate AutoVLA on the nuScenes dataset following both ST-P3 [14] and UniAD [65] protocols, in comparison with state-of-the-art end-to-end models. As shown in Table S2, our model demonstrated competitive performance in the nuScenes planning benchmark.
Fig. S6 shows AutoVLA’s planning outputs. The model generates safe trajectories that closely align with ground-truth motions, accompanied by coherent and context-aware reasoning outputs. However, it can be observed that many nuScenes scenarios are relatively straightforward, often not requiring complex reasoning. This may explain the lack of performance gain in quantitative metrics when reasoning is introduced.
@@
We present additional visualization results of our model on the nuPlan dataset in Fig. S7. In relatively simple scenarios such as curved roads and intersections, our model generates high-quality trajectories via fast thinking. In more complex scenes with numerous traffic regulations, it leverages slow thinking to produce better and regulation-compliant planning results with CoT reasoning. As illustrated in Fig. S8, our AutoVLA model preserves reasoning capabilities in complex scenarios despite removing redundant reasoning after RFT. Qualitative comparisons demonstrate that RFT enhances planning performance in complex scenarios.
@@
Method Name RFS (Overall) $\uparrow$ ADE at 5s (Overall) $\downarrow$ ADE at 3s (Overall)$\downarrow$ RFS (Spotlight)$\uparrow$
@@ F.3 Waymo Results
As of May 22, 2025, AutoVLA achieves competitive performance on the Waymo end-to-end driving Challenge leaderboard, as shown in Table S3. The model ranks highly in both RFS Overall and ADE metrics and achieves the top score in the RFS Spotlight metric, which focuses on the most challenging scenarios. Qualitative results are shown in Fig. S9, where AutoVLA demonstrates its ability to generate safe, context-aware trajectories in complex environments. The model effectively handles interactions, construction zones, and traffic regulations in diverse scenarios while providing coherent reasoning to justify its decisions.
Table S4 presents the ablation studies of the AutoVLA model in different training setups. The results indicate that multi-camera input consistently enhances driving performance. When trained solely on the Waymo end-to-end driving dataset, incorporating reasoning significantly improves performance compared to action-only setups. Moreover, pretraining on nuX data provides a substantial performance boost, suggesting that such pretraining enhances the scene understanding of the model with more driving data. RFT can significantly improve planning performance by mitigating error accumulation in generation and aligning with the task-specific rewards.
@@ F.4 CARLA Results
Closed-loop testing results in the CARLA simulator are shown in Fig. S10. Two representative scenarios are illustrated: (1) the ego vehicle equipped with AutoVLA successfully responds to a cut-in vehicle, and (2) it executes a smooth left turn. Additional closed-loop simulation results are available on the project website.

### v3

@@ E.2 Evaluation Metrics
nuPlan (NAVSIM). PDMS is the official benchmark metric in NAVSIM on the nuPlan dataset, with its formulation provided in Eq.˜S3.
@@
CARLA. We evaluate driving performance in closed-loop testing using four key metrics: Driving Score, Success Rate, Efficiency, and Comfortness following the Bench2Drive benchmark [dauner2024navsim]. Driving Score reflects the overall route completion penalized by infraction severity. Success Rate measures the percentage of routes completed without major infractions within the time limit. Efficiency quantifies how well the ego vehicle maintains a reasonable speed relative to surrounding traffic, computed as the average speed ratio over multiple checkpoints. Comfortness assesses the smoothness of the trajectory based on physical signals such as acceleration, jerk, and yaw rate, using expert-derived thresholds.
@@ E.3 Model Training
Our primary base model is trained on the mixed training set of the nuPlan and nuScenes datasets. The ego state contains the vehicle’s current velocity and acceleration. For the test on the Waymo end-to-end driving dataset, we further fine-tune this base model using the Waymo end-to-end driving dataset. For the model trained on Waymo, the ego vehicle’s state encompasses current acceleration and a 4-second history of vehicle positions and velocities. For closed-loop simulation testing, we train a separate model using data preprocessed from the CARLA Garage dataset and DriveLM-CARLA annotations. This model is trained with single-view inputs instead of the multi-view setup used in the primary model, and we use a large resolution for the input images (with $28\times 28\times 384$ pixels).
@@ E.4 Inference
We employ a stochastic generation strategy using top-p and top-k sampling to generate reasoning and planning outputs. Higher sampling diversity (e.g., temperature=$1.0$, top-p=$0.5$, top-k=2020) supports slow thinking modes, enabling the model to produce deeper and more elaborate reasoning chains. In contrast, more deterministic settings (e.g., temperature=$0.1$, top-p=$0.01$, top-k=11) produce fast thinking, yielding consistent and direct responses.
@@
ST-P3 [hu2022st] 1.33 2.11 2.90 2.11 0.23 0.62 1.27 0.71 - - - - - - - -
VAD [jiang2023vad] 0.17 0.34 0.60 0.37 0.07 0.10 0.24 0.14 - - - - - - - -
UniAD [hu2023planning] 0.44 0.67 0.96 0.69 0.04 0.08 0.23 0.12 0.48 0.96 1.65 1.03 0.05 0.17 0.71 0.31
EMMA [hwang2024emma] 0.14 0.29 0.54 0.32 - - - - - - - - - - - -
OpenEMMA [xing2025openemma] 1.45 3.21 3.76 2.81 - - - - - - - - - - - -
OpenDriveVLA-3B [zhou2025opendrivevla] 0.14 0.30 0.55 0.33 0.02 0.07 0.22 0.10 0.19 0.58 1.24 0.67 0.02 0.18 0.70 0.30
AutoVLA (action only) 0.22 0.39 0.61 0.41 0.10 0.17 0.28 0.18 0.29 0.67 1.17 0.71 0.15 0.34 0.56 0.35
AutoVLA (w/ CoT) 0.21 0.38 0.60 0.40 0.13 0.18 0.28 0.20 0.28 0.66 1.16 0.70 0.14 0.25 0.53 0.31
@@ Appendix F Additional Results
We evaluate AutoVLA on the nuScenes dataset following both ST-P3 [hu2022st] and UniAD [hu2023planning] protocols, in comparison with state-of-the-art end-to-end models. As shown in Table˜S2, our model demonstrated competitive performance in the nuScenes planning benchmark.
Fig.˜S6 shows AutoVLA’s planning outputs. The model generates safe trajectories that closely align with ground-truth motions, accompanied by coherent and context-aware reasoning outputs. However, it can be observed that many nuScenes scenarios are relatively straightforward, often not requiring complex reasoning. This may explain the lack of performance gain in quantitative metrics when reasoning is introduced.
@@
We present additional visualization results of our model on the nuPlan dataset in Fig.˜S7. In relatively simple scenarios such as curved roads and intersections, our model generates high-quality trajectories via fast thinking. In more complex scenes with numerous traffic regulations, it leverages slow thinking to produce better and regulation-compliant planning results with CoT reasoning. As illustrated in Fig.˜S8, our AutoVLA model preserves reasoning capabilities in complex scenarios despite removing redundant reasoning after RFT. Qualitative comparisons demonstrate that RFT enhances planning performance in complex scenarios.
@@
[…]
@@ F.3 Waymo Results
As of May 22, 2025, AutoVLA achieves competitive performance on the Waymo End-to-End Driving Challenge leaderboard, as shown in Table˜S3. The model ranks highly in both RFS Overall and ADE metrics and achieves the top score in the RFS Spotlight metric, which focuses on the most challenging scenarios. Qualitative results are shown in Fig.˜S9, where AutoVLA demonstrates its ability to generate safe, context-aware trajectories in complex environments. The model effectively handles interactions, construction zones, and traffic regulations in diverse scenarios while providing coherent reasoning to justify its decisions.
Table˜S4 presents the ablation studies of the AutoVLA model in different training setups. The results indicate that multi-camera input consistently enhances driving performance. When trained solely on the Waymo end-to-end driving dataset, incorporating reasoning significantly improves performance compared to action-only setups. Moreover, pretraining on nuX data provides a substantial performance boost, suggesting that such pretraining enhances the scene understanding of the model with more driving data. RFT can significantly improve planning performance by mitigating error accumulation in generation and aligning with the task-specific rewards.
@@ F.4 CARLA Results
Closed-loop testing results in the CARLA simulator are shown in Fig.˜S10. Two representative scenarios are illustrated: (1) the ego vehicle equipped with AutoVLA successfully responds to a cut-in vehicle, and (2) it executes a smooth left turn. Additional closed-loop simulation results are available on the project website.
@@
Appendix G Broader Impacts
Autonomous driving is a safety-critical system, further emphasized by the integration of language guidance into the VLA model. This integration necessitates robust safeguards against adversarial attacks and proactive identification and filtering of unsafe human instructions. To mitigate potential hacking threats, it is essential to establish a secure communication channel in the vehicle, complemented by a gated-release strategy for model updates rather than online continual reinforcement fine-tuning on individual vehicles.